OpenStudy (anonymous):

Solve for x withouth using logarithms 2/3= (81/16)^3x

OpenStudy (anonymous):

81 = 3^4 16 = 2^4 hope this helps

OpenStudy (anonymous):

2/3=(16/18)^-3x =(2^4/3^4)^-3x =(2/3)^-12x since bases are same. 1=-12x x=-1/12 answer

OpenStudy (anonymous):

32^x= 4^x+1 Solve without using logarithms

OpenStudy (agent0smith):

\[32^{x}= 4^{(x+1)}\] 32 can be written as 2^5, and 4 as 2^2 \[2^{5x}= 2^{2(x+1)}\] Now can you solve for x?